OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. x + y = 11 -x = -y - 9 A. (10, 1) B. (9, 2) C. (8, 3) D. (11, 0)

OpenStudy (anonymous):

So if -x=-y-9 what does x= ?

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

No. You need to rewrite the -x = -y -9 as x = ?????

OpenStudy (anonymous):

x=y+9?

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

@Godlovesme could you help?

OpenStudy (godlovesme):

yes give me a sec

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

from this x+y=11 solve for x or y

OpenStudy (godlovesme):

@Tallan u there?

OpenStudy (anonymous):

i am one sec

OpenStudy (godlovesme):

k :) jst making sure

OpenStudy (anonymous):

Yea sorry for the short wait I had to record something.. Ok i am confused on how to find it.

OpenStudy (godlovesme):

its fine :) x+y=11 x=11-y got it?

OpenStudy (anonymous):

x=11-y i got that yes.

OpenStudy (godlovesme):

plug "11-y" for x in one of the eqns and solve for y

OpenStudy (anonymous):

0? @Godlovesme

OpenStudy (godlovesme):

can u show me how u got that?

OpenStudy (anonymous):

you plug in x with 11 so it would be 11 +y=11 and it would be 0 to stay at 11

OpenStudy (anonymous):

did i do it right?

OpenStudy (godlovesme):

plug x=11-y in the the -x=-y-9 equation @Tallan

OpenStudy (anonymous):

so it would be -x=-11-9?

OpenStudy (godlovesme):

yes then -(11-y)=-y-9 can u try solving from there?

OpenStudy (anonymous):

It would be 1? y=1

OpenStudy (godlovesme):

yep plug 1 for x in any of the eqns and solve for x

OpenStudy (godlovesme):

i mean plug for y and solve for x :/

OpenStudy (anonymous):

so y=1 so plug 1 in for y?

OpenStudy (godlovesme):

yes sorry for the confusion =0

OpenStudy (godlovesme):

remember x=11-y :)

OpenStudy (anonymous):

10 ?

OpenStudy (godlovesme):

awesome!!!

OpenStudy (godlovesme):

nice job :)

OpenStudy (anonymous):

A would be the answer

OpenStudy (godlovesme):

yep

OpenStudy (godlovesme):

thanks for the medal :)

OpenStudy (anonymous):

np! thanks for the help!!